OpenStudy (anonymous):

Divide 21x^6-6x^4 / 3x^3 simplify your answer as much as possible

OpenStudy (anonymous):

@sleepyhead314

OpenStudy (sleepyhead314):

try factoring the top first :)

OpenStudy (anonymous):

do I multiply 21(6) or subtract and what do I do with the exponents?

OpenStudy (sleepyhead314):

hehe just move the Greatest Common Factor out to the front for example if I had 24y^7 - 16y^5 it would become 8y^5 (3y^2 - 2) can you try that with your problem? :)

OpenStudy (sleepyhead314):

hint: the GCF is 3x^4

OpenStudy (anonymous):

3x^4(2x^2)?

OpenStudy (sleepyhead314):

mmm close! 3x^4 (7x^2 - 2) do you kinda get where I got that or would you like for me to break it down for you more? :)

OpenStudy (anonymous):

can you break it down more?

OpenStudy (sleepyhead314):

ok! we have 21x^6 - 6x^4 on top 21x^6 is basically the multiplied version of (3x^4)*(7x^2) and 6x^4 is the multiplied version of (3x^4)*(2) do you follow? :)

OpenStudy (anonymous):

yes

OpenStudy (sleepyhead314):

then like how ab - ac = a (b - c) 21x^6 - 6x^4 = (3x^4)*(7x^2) - (3x^4)*(2) = (3x^4) (7x^2 - 2)

OpenStudy (anonymous):

ok so the top is (3x^4)(7x^2-2) ?

OpenStudy (sleepyhead314):

yep! :)

OpenStudy (sleepyhead314):

now, what would you get if it was just 3x^4 / 3x^3 ?

OpenStudy (anonymous):

1x^7

OpenStudy (sleepyhead314):

try again ^_^ when it is dividing, the exponents Subtract :)

OpenStudy (anonymous):

1x?

OpenStudy (sleepyhead314):

yep!!! hehehe so you can now get x(7x^4 - 2) :P

OpenStudy (sleepyhead314):

oh wait, I meant x(7x^2 - 2)

OpenStudy (anonymous):

is that the answer? im sorry lol

OpenStudy (sleepyhead314):

all you have to do is distribute ^_^

OpenStudy (sleepyhead314):

you don't have to apologize ;) I was explaining it a little weird xD

OpenStudy (anonymous):

so 7x^3-2x?

OpenStudy (sleepyhead314):

yep, that would be your answer :D